OpenStudy (anonymous):

Part 1: Show the work to solve |5x – 2| ≥ 8. Part 2: Describe the graph of the solution in words.

hero (hero):

1. On a hidden sheet of paper (and while no one is looking) perform the following step: \( -8 \ge 5x - 2 \ge 8\) 2. Then on the actual piece of paper, re-write the above as two separate inequalities: \(5x - 2 \ge 8\) \(5x - 2 \le -8\) 3. Continue solving from there.

OpenStudy (anonymous):

more please! lol I dont understand

OpenStudy (anonymous):

I need a step by step answer ;/